OpenStudy (anonymous):

Find the equations of the tangent and normal to the given curves at the indicated points:

OpenStudy (anonymous):

\[y = x^4 - 6x^3 + 13x^2 - 10x + 5\] at (0,5) \[y = x^4 - 6x^3 + 13x^2 - 10x + 5\] at (1,3) \[y = x^3 at (1,1)\]

OpenStudy (imstuck):

The derivative of the first equation is\[y \prime=4x ^{3}-18x ^{2}+26x-10\]At the point (0,5), the slope is -10.

OpenStudy (imstuck):

The equation would use the point (0,5) to make: y-5=-10(x-0) y=-10x+5 So the normal line's equation would be y = 1/10x + 5

OpenStudy (imstuck):

At the point (1,3), the slope of the tangent line is 2, and the equation would be y-3=2(x-1) y-3=2x-2 y=2x+1 So the equation for the normal line would be y = -1/2x + 1

OpenStudy (imstuck):

The last one has a derivative of \[y=3x ^{2}\]so at (1,1) the slope is 3. The equation would be y-1=3(x-1) y-1=3x-3 y=3x-2 So the equation for the normal line is y=-1/3x-3

OpenStudy (imstuck):

Are you with me on this?

OpenStudy (anonymous):

I'm so sorry, hang on ill do this on paper

OpenStudy (anonymous):

So we use (0, 5) with 0 for x and 5 for y?